OpenStudy (anonymous):

write an equation of the line containing the given point and parallel to the given line. (2,-6) 5x-8y=3

OpenStudy (anonymous):

Okay: First we put the equation, given in standard form in y=mx+b format. 5x-8y=3 subtract 5x -8y=-5x+3 divide by -8 y=(-5/8)x-(3/8) -5/8 is your slope (it is in the m position for slope) now you have to put your point and slope in an equation called point-slope form: \[y-y _{1}=m(x-x _{1})\] With Numbers: \[y-(-6)=(-5/8)(x-2)\] simplify this to: y+6=(-5/8)x+1.25 This finnally leads to the final answer: y=(-5/8x)-(19/4)